write the equation of the best-fit line in slope intercept form using the two data points given (9,2)(6,9)
Answer by stanbon(75887)

write the equation of the best-fit line in slope intercept form using the two data points given (9,2)(6,9)
----------------------
slope = (9-2)/(6-9) = 7/-3 = -7/3
-----
Form y = mx+b
2 = (-7/3)9 + b
2 = -21 + b
b = 23
--------
Equation:
y = (-7/3)x+23
======================
Cheers,
Stan H.